OpenStudy (anonymous):

okay guys help me one this one -x+5y=1 2x+4y=-4

OpenStudy (saifoo.khan):

\[\Large -\frac{12}{7}, -\frac{1}{7}\]

OpenStudy (anonymous):

I'm glad laura has learned from your help, Saifoo. Now she's guaranteed to make an A on the test. >_>

OpenStudy (anonymous):

that is def not the anwser ha .. wrong form .. you are suppose to do the same thing to this as my last post

OpenStudy (saifoo.khan):

lol, this is the correct answer untill your equation is incorrect.

OpenStudy (anonymous):

no my equation is not wrong

OpenStudy (saifoo.khan):

Then my answer isnt too!

OpenStudy (anonymous):

yes your answer is wrong, i don't think you know what your doing exactly no offense

OpenStudy (saifoo.khan):

lol. srsly my answer isnt incorrect.

OpenStudy (saifoo.khan):

Let me prove it.

OpenStudy (anonymous):

okayy well none of the answers on the page have what you put for the answer

OpenStudy (anonymous):

saifoo's answer is correct x=-12/7 y=-1/7

OpenStudy (saifoo.khan):

\[\huge -(-\frac{12}{7})+ 5(-\frac{1}{7}) = 1 \]

OpenStudy (saifoo.khan):

thnk Pro. :]

OpenStudy (anonymous):

okayy thanks

OpenStudy (anonymous):

sorry i was looking at the wrong problem

OpenStudy (saifoo.khan):

Lol.